# Vamos: Versatile Action Models for Video Understanding

### v1

@@ Abstract
What makes good video representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as discrete action labels, or free-form video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularities. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embeddings, action labels, and free-form descriptions extracted from videos as its input. We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, Next-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We perform extensive ablation study and qualitative analysis to support our observations, and achieve state-of-the-art performance on three benchmarks. Code and models will be released at https://brown-palm.github.io/Vamos/.
@@ 1 Introduction
Figure 1: Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations, discrete action labels, and free-form text descriptions for downstream applications.
A picture held us captive … for it lay in our language and language seemed to repeat it to us inexorably.
Philosophical Investigations
@@
We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [41, 48] on sampled video frames, serve as useful video representations for action modeling from videos? And if so, can we then leverage a pre-trained large language model (LLM) [64] as the generative model of actions, represented as free-form text? Despite the apparent crudeness, we argue that it is inspired and supported by recent advances in LLM research, where LLMs have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [22, 52], and performing reasoning [4, 13, 70, 14], all of which are indispensable for action modeling.
@@
To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings, discrete action labels, and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [64]. The visual embeddings are linearly projected into the same language space following standard practice [41, 53]. As illustrated in Figure 1, the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type. Vamos can directly leverage an LLM’s next token prediction capability for action anticipation [21]. Similar to how an LLM can be tasked to perform reasoning, we ask Vamos to perform video question answering [71], by prepending the question to the video representation.
@@
We perform extensive evaluations on four benchmarks, including the Ego4D [21] long-term action anticipation, NeXT-QA [71] and IntentQA [43] video question answering, and EgoSchema [50] zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the representation based on discrete action labels outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [68] by 50%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings only result in marginal performance gains.
Another benefit of text-based video representation is interpretability. We exploit this property in two scenarios: First, we perform causal intervention and correct the suboptimal video descriptions, and we observe that the answers are indeed corrected accordingly. Second, we propose a lightweight token selector module to sample as few as 20 tokens from video descriptions that are more than 600 tokens long. We observe that not only the video question answering performance remains comeptitive, outperforming the visual embedding baseline, but the selected tokens are indeed highly related to answering the questions, confirming that the Vamos framework behaves as intended.
To summarize, we make the following contributions:
  1. 1.
    We propose to revisit video representation and action modeling in the LLM era, and to explore the effectiveness of free-form text descriptions.
  2. 2.
    We propose Vamos, a versatile framework that allows us to incorporate and compare the effectiveness of different video representations, both visual and textual.
  3. 3.
    We demonstrate the effectiveness of free-form text as a representation that is performant, interpretable, and can be intervened, even when aggressively compressed to 6%. Vamos achieves state-of-the-art performance on Ego4D LTA, NeXT-QA, IntentQA, and EgoSchema.
@@ 2 Related Work
Figure 2: Overview of Vamos and token selector. (a) Vamos flexibly unifies distributed visual features and textual video representations including discrete action labels and free-form video captions. By performing sequence modeling, Vamos can be applied to multiple video understanding tasks. (b) The lightweight token selector as an add-on module of Vamos. It pick one single token from the video sequence based on the task sequence. By applying the token selector to $k$ segments of textual video representation, $k$ tokens are finally selected as the compact input to the LLM for down-stream tasks.
@@
Vision-Language Foundation Models. Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension. Recent advancements in vision-language models have been propelled by various pretraining techniques leveraging large-scale multimodal datasets. Models such as CLIP [58], ALIGN [33], CLOOB [19] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs. Another line of vision-language models showcased by VisualBERT [44], UNITER [11], OSCAR [46], FLAVA [61], uses a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection. For videos, VIOLET [17] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [62, 17, 81, 82]. The objectives can be combined [78] and the encoders for different modalities can be shared [66]. Compared to existing vision-language models, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
@@
Visually-augmented LLMs. Apart from joint visual-language pre-training, existing large language models (LLMs) can also be augmented to incorporate visual inputs. For example, VisualGPT [9] and Flamingo [2] directly fuse visual information into the layers of a language model decoder using a cross-attention mechanism instead of using images as additional prefixes to the language model. Other approaches, such as instructional tuning [48], or linearly projecting the visual embeddings into the input space of LLMs [41, 53], have also been explored. Vamos largely follows this linear projection approach to incorporate visual embedding inputs, with the goal to understand if and how they are complementary to text-based video representations.
@@
Additionally, tool-using large language models have been recently proposed to invoke and incorporate the use of task-specific modules [60, 27], where visual perceptions consist a substantial subset of the tools. Notably, VisualProgram [24] and ViperGPT [63] propose to apply LLMs to generate symbolic programs based on pre-selected computer vision modules for visual question answering. Closest to our work is Socratic Models [83], where the authors propose to use natural language as the common interface to connect foundation models with different input and output modalities. Vamos differs from these approaches as it aims to understand the power and limitations of text-based representation for a wide range of video understanding tasks, through quantitative experiments and qualitative analysis.
@@ 3.1 Text-based Video Representation
Compared with a single image, the video data often contains more complex and dynamic information including context and interactions. Prior works [39, 80, 69] have demonstrated the effectiveness of condensing images into conceptual or natural language bottleneck with good interpretability, while it is still unclear if long form videos with richer dynamics can also be condensed into such explainable textual representations. To translate videos into language bottleneck, we consider two methods: discrete action labels and free-form text descriptions.
We propose to use an extremely compact representation by condensing videos into sequences of discrete action labels, offering a more compact representation. In practice, this is achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
Video text descriptions use free-form natural language to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video. It transforms complex and dynamic visual information into more comprehensible and interpretable textual descriptions. We employ off-the-shelf image captioning models such as BLIP-2 [42] that process sampled frames to produce frame-level captions, which are subsequently concatenated to form a comprehensive video-level caption.
@@ 3.2 Versatile Action Models
Benefiting from the encoded extensive prior knowledge, large language models (LLMs) have demonstrated strong capability for temporal and causal reasoning, which are crucial for video understanding. We introduce Vamos, a simple yet effective framework to utilize LLMs to unify video dynamic modeling tasks, including comprehending historical content (video question answering, VQA) and future prediction (long-term action anticipation, LTA). As shown in Figure 2 (a), given a video $V$ and a pretrained LLM $\pi$, the input sequence $x_{t}=[x_{tvr},x_{task}]$ consists of the textual video representations $x_{tvr}$ of $V$ and other task specific langauge inputs $x_{task}$ (e.g., instructions, questions, targets). The frozen word embedding layer $\mathcal{F}_{emb}$ first generate the corresponding text tokens $z_{t}=\mathcal{F}_{emb}(x_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $x_{t}$, $D$ is the feature dimension.
@@
To address any residual information not entirely captured by the text-based representations, we incorporate a learnable linear projection layer $\mathcal{F}_{proj}$ to align visual features with the language space. Specifically, the frozen vision backbone $\mathcal{E}$ takes in $N_{v}$ frames $[v_{1},...v_{N_{v}}]$ sampled from $V$ to generate the visual features. These visual features are then input into the projection layer $\mathcal{F}_{proj}$ to produce vision tokens $z_{v}=\mathcal{F}_{proj}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$. $z_{v}$ and $z_{t}$ are then concatenated and fed into the LLM $\pi$. During training, the LLM $\pi$ is optimized with the standard language modeling loss using LoRA [26].
@@
Vamos is accommodating multiple video understanding tasks that can be formatted as sequence modeling tasks with varying task-specific inputs $x_{task}$ and modeling objectives. Specifically, for the VQA task, $x_{task}$ is composed of instructions, questions, and answers, with the answer being the training objective. During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA, depending on the dataset. For the LTA task, $x_{task}$ is composed of instructions and future actions, where the training objective is the future action sequence. During inference, the trained LLM generates future action sequences based on history actions.
3.3 Lightweight Token Selector
We also investigate the potential of compressing long input sequences and extracting keys element from the generated textual representations to probe the crucial information for downstream video understanding tasks. We design a lightweight token selector as a add-on module (shown in Figure 2 (b)) for Vamos. The token selector takes in a sequence of textual video tokens to be selected, and a task-specific token sequence, functioning as queries. It picks a single token from the video tokens that is most relevant for downstream video understanding. To select a condensed token sequence for a video, the original textual video token sequence with $L$ tokens is first uniformly divided into $k\ll L$ segments $\{X^{(1)}$,…, $X^{(k)}\}$ containing $n$ tokens. Each segment $X^{(i)}=\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ is fed into the token selector as candidate tokens, from which one token $x^{(i)}$ is selected based on the task texts (question for VQA).

### v2

@@ Abstract
What makes good representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as general-purpose video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularity. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embedding and free-form text descriptions as its input. To interpret the important text evidence for question answering, we generalize the concept bottleneck model to work with tokens and nonlinear models, which uses hard attention to select a small subset of tokens from the free-form text as inputs to the LLM reasoner. We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, NeXT-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We also demonstrate that our token bottleneck model is able to select relevant evidence from free-form text, support test-time intervention, and achieves nearly 5 times inference speedup while keeping a competitive question answering performance. Code and models are publicly released at https://brown-palm.github.io/Vamos/.
@@ 1 Introduction
Figure 1: (a) Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations and free-form text descriptions for downstream applications. (b) Overview of the Vamos framework. It flexibly unifies distributed visual features and textual video representations such as free-form video captions. By performing sequence modeling, Vamos can be applied to multiple video understanding tasks.
@@
We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [40, 48] on sampled video frames, serve as useful video representations for action modeling from videos? And if so, can we then leverage a pre-trained large language model (LLM) [67] as the generative model of actions, represented as free-form text? If both answers are yes, visual reasoning can then enjoy recent advances in LLM research, where they have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [19, 53], and performing reasoning [4, 12, 74, 13], all of which are indispensable for action modeling. Since we assume the text descriptions are general purpose, they can be extracted once and reused for different downstream tasks, similar to pre-computed visual embeddings [59].
@@
To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [67]. The visual embeddings are linearly projected into the same language space following standard practice [40, 54]. As illustrated in Figure 1(a), the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type. Vamos can directly leverage an LLM’s next token prediction capability for action anticipation [18]. We also ask Vamos to perform video question answering [75], by appending the question to the video representation as inputs to the LLM reasoner.
One inherent benefit of text-based video representation is its interpretability. Inspired by interpretable object classifiers such as concept bottleneck models (CBM) [37], we aim to understand which words serve as important evidence for question answering. However, CBM requires a pre-defined list of discrete visual concepts, and requires a linear classifier to achieve interpretability. We generalize this framework and learns hard attention to select a small subset of text inputs to feed to the LLM reasoner, where the text inputs are tokenized text as opposed to pre-defined concepts. We call our generalized formulation token bottleneck models (TBM). TBM naturally supports the incorporation of multimodal information, and allows users to perform causal intervention.
@@
We perform extensive evaluations on four benchmarks, including the Ego4D dataset [18] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, and EgoSchema [50] for zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the text-based representation outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 50%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings surprisingly results in marginal or no performance gains. Finally, we demonstrate that the token bottleneck model is able to select semantically relevant evidence for question answering, and achieves 5x speedup at inference time while maintaining the question answering accuracy.
@@ 2 Related Work
Vision-Language Foundation Models. Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension. Models such as CLIP [59], ALIGN [30], CLOOB [17] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs. Another line of vision-language models [43, 10, 45, 63] use a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection. For videos, VIOLET [16] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [64, 16, 86, 85]. The objectives can be combined [82] and the encoders for different modalities can be shared [69]. Compared to existing VLMs, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
@@
Visually-augmented LLMs. Apart from joint visual-language pre-training, existing large language models (LLMs) can also be augmented to incorporate visual inputs. For example, VisualGPT [9] and Flamingo [2] directly fuse visual information into the layers of a language model decoder using a cross-attention mechanism instead of using images as additional prefixes to the language model. Other approaches, such as instructional tuning [48], prompting large language models for knowledge retrieval [62], or linearly projecting the visual embeddings into the input space of LLMs [54, 46], have also been explored. Vamos largely follows this approach to incorporate visual embeddings, with the goal to understand if and how they are complementary to text-based video representations.
@@
Additionally, tool-using large language models have been recently proposed to invoke and incorporate the use of task-specific modules [61, 24], where visual perceptions consist a substantial subset of the tools. Notably, VisualProgram [21] and ViperGPT [65] propose to apply LLMs to generate symbolic programs based on pre-selected computer vision modules for visual question answering. VidIL [72] leverages expert knowledge to design object and action concepts for few-shot captioning and video question answering. Closest to our work is Socratic Models [87], where the authors propose to use natural language as the common interface to connect foundation models with different input and output modalities. Finally, several concurrent works [88, 42, 51] share similar motivations and methods to Vamos in utilizing text-based representations and modules for video understanding, without the incorporation of visual embeddings.
@@ 3.1 Text-based Video Representation
A video often contains complex and dynamic information including context and interactions. While prior works [37, 84, 73] have demonstrated the effectiveness of condensing images into text-based representations such as visual concepts, it remains unclear if videos can also be condensed into text-based representations. To answer this research question, we consider text descriptions that are task-agnostic, and can potentially be applied in diverse video understanding tasks.
Concretely, we rely on general-purpose captioning models to generate free-form text descriptions to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video. We employ off-the-shelf image captioning models such as BLIP-2 [41] that generate image-level captions from the sampled video frames. These captions are subsequently concatenated to form a comprehensive video-level caption.
For certain tasks, prior knowledge might be helpful to guide the model learning. For example, when the goal is to model the long-term temporal dynamics of verbs and nouns for the long-term action anticipation task, it would be beneficial to trim the inputs to only contain discrete action labels. In practice, this can be achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
@@ 3.2 Versatile Action Models
Large language models (LLMs) have demonstrated strong capability for temporal reasoning [90] and even some potential for causal reasoning [35], both of which are crucial for video understanding. We introduce Vamos, a simple yet effective framework to utilize LLMs to unify video dynamic modeling tasks, including comprehending historical content (video question answering, VQA) and future prediction (long-term action anticipation, LTA). As shown in Figure 1 (b), given a video $V$ and a pretrained LLM $\pi$, the input sequence $\mathbf{x}_{t}=[\mathbf{x}_{\text{tvr}},\mathbf{x}_{\text{task}}]$ consists of the textual video representations $\mathbf{x}_{\text{tvr}}$ of $V$ and other task specific language inputs $\mathbf{x}_{\text{task}}$ (e.g., instructions, questions, targets). The frozen word embedding layer $\mathcal{F}_{\text{emb}}$ first generate the corresponding text tokens $\mathbf{z}_{t}=\mathcal{F}_{\text{emb}}(\mathbf{x}_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $\mathbf{x}_{t}$, $D$ is the feature dimension.
@@
Vamos incorporates the residual information not entirely captured by $\mathbf{x}_{\text{tvr}}$ via representations encoded directly from the visual modality, such as CLIP visual embedding. We adopt a learnable linear projection layer $\mathcal{F}_{\text{proj}}$ to align visual features with the language space. Specifically, the frozen vision backbone $\mathcal{E}$ takes in $N_{v}$ frames $[v_{1},...v_{N_{v}}]$ sampled from $V$ to generate the visual features. These visual features are then fed into the projection layer $\mathcal{F}_{\text{proj}}$ to produce visual tokens $\mathbf{z}_{v}=\mathcal{F}_{\text{proj}}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$. To combine information from the visual and textual representations, we adopt the early fusion strategy and concatenate $\mathbf{z}_{v}$ and $\mathbf{z}_{t}$ as the inputs to the LLM $\pi$. When labeled training data is available for task-specific fine-tuning, we update the weights of the LLM $\pi$ either with LoRA [23] or LLaMA-Adapter [89].
@@
Vamos can accommodate diverse video understanding tasks by formulating each task as sequence completion given an appropriate task description $\mathbf{x}_{\text{task}}$, the LLM $\pi$ can then be optimized with the standard language modeling objectives. Specifically, for the VQA task, $\mathbf{x}_{\text{task}}$ is composed of instructions, questions, and answers, with the answer being the training objective. During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA. For the LTA task, $\mathbf{x}_{\text{task}}$ is composed of instructions and future actions, where the training objective is the future action sequence. During inference, the fine-tuned LLM is tasked to generate sequences of future actions based on the history actions.
Discussion: By design, Vamos naturally incorporates temporal information as the captions are timestamped. Since the text-based representation is general-purpose, the framework is efficient since once extracted, the text-based representation can be reused for different questions, just like the CLIP embeddings commonly used by VLMs. Vamos is a two-stage framework with decoupled “perception” and “reasoning” modules, in addition to the benefits on interpretability (of the intermediate representation) and efficiency (reuse of the intermediate features). Another conceptual benefit is its generalizability: While the visual distributions of different datasets may differ, the reasoning module may be shared. We show this is indeed the case for EgoSchema [50], when the end-to-end vision-language models performance significantly worse than Vamos.
3.3 Token Bottleneck Models
Figure 2: An illustration of the token bottleneck model (TBM). We are inspired by the concept bottleneck models (CBM) [37], which achieve interpretable object classification by inspecting the weights of the learned linear classifier (left). Unlike CBM, Vamos does not require pre-defining a list of concepts. It directly works with tokenized text inputs. To provide input tokens to the reasoning model (an LLM), we leverage hard attention to generate binary rather than continuous weights (middle). The token bottleneck can be interpreted directly. It can also be intervened with human inputs (right), or augmented with residual visual information.
We aim to understand how the text-based representations are utilized by the LLM reasoners. As an LLM operates like a black box, we aim to enhance the interpretability of the overall framework by first understanding how it selects evidence to solve the downstream tasks. As illustrated in Figure 2, we are inspired by the success of interpretable object classifiers, such as the concept bottleneck model (CBM) [37]. CBM relies on pre-defining a list of concepts and building (often supervised) concept detectors for each of them. As our model’s inputs are free-form text, we propose to directly work with word tokens as opposed to pre-defined concepts. In addition, CBM relies on linear classifiers to achieve model interpretability. Each weight in a learned classifier indicates the importance of the corresponding concept for making the prediction. We hypothesize that linear classifiers are not sufficiently expressive when solving tasks that require a stronger “reasoner”. To strike a balance between model interpretability and expressiveness, we generalize the CBM framework to learn binary attention on the input tokens as opposed to continuous weights used by the CBM linear classifier. As illustrated in Figure 2, The binary weights indicate which tokens are to be selected and fed to the more expressive LLM for solving the target tasks. We name this generalized framework as token bottleneck models (TBMs).
To implement TBM, we design a lightweight token selector as an add-on module for Vamos (Figure 2 middle). The token selector takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs. It learns to pick a single token among the candidate tokens $\mathbf{z}_{\text{tvr}}$ by optimizing the objective for a given downstream task. To select a sequence of tokens, we assume the important information is even distributed across the input sequence, and uniformly divide the input sequence into $k$ segments $\{\mathbf{z}_{\text{tvr}}^{(1)}$,…, $\mathbf{z}_{\text{tvr}}^{(k)}\}$, each of which contains $n$ tokens. Each segment $\mathbf{z}_{\text{tvr}}^{(i)}=\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ is fed into the token selector, from which one token $z^{(i)}$ is selected for the task $\mathbf{z}_{\text{task}}$.